# NDKF: A Neural-Enhanced Distributed Kalman Filter for Nonlinear Multi-Sensor Estimation

### v1

@@ III Convergence and Stability Analysis
This section provides a rigorous analysis of the stability of NDKF under local linearization and distributed fusion.
@@ III-A Assumptions
A few assumptions will be used in the analysis:
@@
    Smoothness and Lipschitz Conditions: The neural network functions $f_{\theta}(\cdot)$ and $h_{\theta,i}(\cdot)$, for each node $i$, are continuously differentiable and locally Lipschitz in a neighborhood $\mathcal{X}$ of the true state trajectory. That is, there exist constants $L_{f}>0$ and $L_{h_{i}}>0$ such that for all $\mathbf{x},\mathbf{y}\in\mathcal{X}$,
    $\|f_{\theta}(\mathbf{x})-f_{\theta}(\mathbf{y})\|$ $\leq L_{f}\|\mathbf{x}-\mathbf{y}\|,$ (17)
    $\|h_{\theta,i}(\mathbf{x})-h_{\theta,i}(\mathbf{y})\|$ $\leq L_{h_{i}}\|\mathbf{x}-\mathbf{y}\|.$ (18)
@@
    Bounded Jacobians: The Jacobians
    $\mathbf{F}_{\theta,k}=\nabla f_{\theta}(\hat{\mathbf{x}}_{k|k})\quad\text{and}\quad\mathbf{H}_{\theta,i,k}=\nabla h_{\theta,i}(\hat{\mathbf{x}}_{k|k})$
    are uniformly bounded in norm by constants $M_{f}$ and $M_{h_{i}}$, respectively.
@@
    Noise and Initial Error Bounds: The process noise $\mathbf{w}_{k}$ and the measurement noise $\mathbf{v}_{k,i}$ are bounded in mean-square sense. Moreover, there exists a constant $\delta>0$ such that the initial estimation error satisfies
    $\|\mathbf{e}_{0|0,i}\|<\delta,\quad\text{for all nodes }i.$
@@ III-B Local Error Dynamics
Under the local linearization, the prediction step yields an a priori error dynamics approximated by
$\mathbf{e}_{k+1|k}\approx\mathbf{F}_{\theta,k}\,\mathbf{e}_{k|k,i}+\mathbf{w}_{k},$
while the measurement update gives
$\mathbf{e}_{k|k,i}\approx\bigl{(}\mathbf{I}-\mathbf{K}_{k,i}\,\mathbf{H}_{\theta,i,k}\bigr{)}\,\mathbf{e}_{k|k,i}-\mathbf{K}_{k,i}\,\mathbf{v}_{k,i}.$
The distributed fusion step is assumed to combine local estimates (in the information form) without degrading these contraction properties.
@@ III-C Sufficient Conditions for Stability
The following theorem provides sufficient conditions for the local stability of NDKF.
Theorem 1 (Stability of NDKF)
Under assumptions (i)–(iii), suppose that there exist constants $\alpha,\beta\in(0,1)$ such that for all time indices $k$ and for each node $i$,
$\|\mathbf{F}_{\theta,k}\|\leq\alpha,\quad\text{and}\quad\|\mathbf{I}-\mathbf{K}_{k,i}\,\mathbf{H}_{\theta,i,k}\|\leq\beta.$ (19)
Then, there exists a neighborhood $\mathcal{E}$ around the true state such that if $\|\mathbf{e}_{0|0,i}\|<\delta$, the estimation error $\mathbf{e}_{k|k,i}$ converges exponentially fast (in mean-square sense) to a bounded region determined by the process and measurement noise covariances.
@@
We first analyze the error propagation for a single node $i$ before considering the fusion step. For clarity, let $\mathbf{e}_{k|k,i}=\mathbf{x}_{k}-\hat{\mathbf{x}}_{k|k,i}$ denote the local estimation error at time $k$, and assume that the local linearization is exact in the sense that higher-order terms are negligible for $\mathbf{e}_{k|k,i}$ sufficiently small.
@@
while the predicted state is
@@
By the mean value theorem and (A.i), there exists a point $\xi_{k}$ on the line segment between $\mathbf{x}_{k}$ and $\hat{\mathbf{x}}_{k|k,i}$ such that
@@
where $\mathbf{F}_{\theta}(\xi_{k})$ is the Jacobian evaluated at $\xi_{k}$. Using (A.ii), we have
$\|\mathbf{F}_{\theta}(\xi_{k})\|\leq M_{f}\leq\alpha<1.$ (23)
Thus, the a priori error becomes
$\mathbf{e}_{k+1|k,i}=f_{\theta}(\mathbf{x}_{k})-f_{\theta}(\hat{\mathbf{x}}_{k|k,i})+\mathbf{w}_{k}=\mathbf{F}_{\theta}(\xi_{k})\mathbf{e}_{k|k,i}+\mathbf{w}_{k},$ (24)
and taking the norm yields
$\|\mathbf{e}_{k+1|k,i}\|\leq\|\mathbf{F}_{\theta}(\xi_{k})\|\,\|\mathbf{e}_{k|k,i}\|+\|\mathbf{w}_{k}\|\leq\alpha\,\|\mathbf{e}_{k|k,i}\|+\|\mathbf{w}_{k}\|.$ (25)
ii. Measurement Update. At the measurement update, the local filter uses the measurement
@@
By subtracting the update equation from the true state (20) and using a similar linearization for $h_{\theta,i}(\cdot)$ (with Jacobian $\mathbf{H}_{\theta,i,k+1}$ satisfying $\|\mathbf{H}_{\theta,i,k+1}\|\leq M_{h_{i}}$ by (A.ii)), we obtain
$\mathbf{e}_{k+1|k+1,i}$ $=\mathbf{x}_{k+1}-\hat{\mathbf{x}}_{k+1|k,i}$
$-\mathbf{K}_{k+1,i}\Bigl{(}h_{\theta,i}(\mathbf{x}_{k+1})-h_{\theta,i}(\hat{\mathbf{x}}_{k+1|k,i})+\mathbf{v}_{k+1,i}\Bigr{)}$
$=\Bigl{(}\mathbf{I}-\mathbf{K}_{k+1,i}\mathbf{H}_{\theta,i,k+1}\Bigr{)}\mathbf{e}_{k+1|k,i}-\mathbf{K}_{k+1,i}\mathbf{v}_{k+1,i}.$ (29)
Taking norms and applying the triangle inequality, we have
@@
By condition (19),
$\|\mathbf{I}-\mathbf{K}_{k+1,i}\mathbf{H}_{\theta,i,k+1}\|\leq\beta<1.$ (31)
Combining this with (25) gives
$\|\mathbf{e}_{k+1|k+1,i}\|\leq\beta\Bigl{(}\alpha\,\|\mathbf{e}_{k|k,i}\|+\|\mathbf{w}_{k}\|\Bigr{)}+\|\mathbf{K}_{k+1,i}\|\,\|\mathbf{v}_{k+1,i}\|.$ (32)
Define $\gamma=\alpha\beta$, which is less than 1 since both $\alpha$ and $\beta$ lie in $(0,1)$. Then,
$\|\mathbf{e}_{k+1|k+1,i}\|\leq\gamma\,\|\mathbf{e}_{k|k,i}\|+\beta\,\|\mathbf{w}_{k}\|+\|\mathbf{K}_{k+1,i}\|\,\|\mathbf{v}_{k+1,i}\|.$ (33)
iii. Induction and Contraction. Define the noise term
$\nu_{k}=\beta\,\|\mathbf{w}_{k}\|+\|\mathbf{K}_{k+1,i}\|\,\|\mathbf{v}_{k+1,i}\|.$ (34)
Then, (33) can be written as
$\|\mathbf{e}_{k+1|k+1,i}\|\leq\gamma\,\|\mathbf{e}_{k|k,i}\|+\nu_{k}.$ (35)
Iterating this inequality for $k{=}0$ to $N$, we obtain by induction
$\|\mathbf{e}_{N|N,i}\|\leq\gamma^{N}\|\mathbf{e}_{0|0,i}\|+\sum_{j=0}^{N-1}\gamma^{N-1-j}\nu_{j}.$ (36)
Since $\gamma<1$, the term $\gamma^{N}\|\mathbf{e}_{0|0,i}\|$ decays exponentially fast. Under (A.iii), the noise terms $\nu_{j}$ are bounded in mean-square; hence, the summation converges to a finite value as $N\to\infty$. This implies that $\|\mathbf{e}_{N|N,i}\|$ converges exponentially fast to a neighborhood of the origin whose size is determined by the noise bounds.
iv. Distributed Fusion Step. Assume that after the local update, each node participates in a distributed fusion step where the fused estimate is computed as a convex combination (or via an information form aggregation) of the local estimates. Let $\mathcal{N}_{i}$ denote the set of nodes communicating with node $i$, and let the fused error be
@@
with weights $w_{ij}{\geq}0$ and $\sum_{j\in\mathcal{N}_{i}}w_{ij}{=}1$. Then, by convexity,
$\|\mathbf{e}^{(\mathrm{fusion})}_{k|k,i}\|\leq\sum_{j\in\mathcal{N}_{i}}w_{ij}\,\|\mathbf{e}_{k|k,j}\|\leq\max_{j\in\mathcal{N}_{i}}\|\mathbf{e}_{k|k,j}\|.$ (38)
Thus, the fusion step does not increase the maximum error among the nodes. As a result, the contraction established in (36) for each node is preserved under fusion.
Combining the above results, we conclude that, for each node $i$, there exist constants $C,C^{\prime}>0$ such that
$\mathbb{E}\Bigl{[}\|\mathbf{e}_{k|k,i}\|^{2}\Bigr{]}\leq C\,\gamma^{2k}\|\mathbf{e}_{0|0,i}\|^{2}+\frac{C^{\prime}}{1-\gamma^{2}}.$ (39)
This inequality shows that the estimation error converges exponentially fast (in mean-square) to a bounded region determined by the noise covariances, establishing the local stability of the NDKF under the stated conditions. ∎
@@ IV Results and Discussion
In this section, we present empirical evaluations of the proposed Neural-Enhanced Distributed Kalman Filter (NDKF) on a 2D system monitored by four distributed sensor nodes, and compare its performance to an Extended Kalman Filter (EKF) baseline.
@@ IV-A Experimental Setup
System Simulation: We consider a 2D state vector
$\mathbf{x}_{k}=[p_{x},\,p_{y}]^{T},$
which evolves according to
$\mathbf{x}_{k+1}\;=\;\mathbf{x}_{k}+\begin{bmatrix}0.05\cos\!\bigl{(}k/10\bigr{)}\\[4.0pt] 0.05\sin\!\bigl{(}k/10\bigr{)}\end{bmatrix}+\boldsymbol{\omega}_{k},$ (40)
@@
Neural Network Configuration: The dynamics residual network in the NDKF is designed to learn the correction to the nominal dynamics. Its input is the state augmented with a time parameter, and it is implemented as a deep fully connected network with three hidden layers (128 neurons each), batch normalization, and dropout (0.2). It is trained for 3000 epochs using Adam with an initial learning rate of 0.001 and a learning rate scheduler. The measurement networks (one per sensor) are implemented with two hidden layers (32 neurons each) and are trained for 1000 epochs.
@@ IV-B Experimental Evaluation
Figure 1 shows a representative 2D trajectory comparing the true state, the local node estimates, and the fused state estimate obtained using the NDKF. The NDKF’s fused estimate closely tracks the true state, while the EKF baseline deviates considerably, as indicated by the numerical RMSE in Table I, due to its mis‑specified measurement models. Figure 2 illustrates the measurement innovation residuals for Node 1. The low and stable residuals obtained with the NDKF, along with the numerical results reported in Table I, confirm its superior measurement update performance compared to the EKF baseline.
@@
An implementation of NDKF, including the example provided in this section, is available in the GitHub repository accompanying this paper at:
https://github.com/sfarzan/NDKF

### v2

@@ III Convergence and Stability Analysis
This section provides a rigorous analysis of the stability of the NDKF under local linearization and distributed fusion.
@@ III-A Assumptions
The following assumptions are used in the analysis:
@@
    Local smoothness on a compact operating set: There exists a compact set $\mathcal{X}\subset\mathbb{R}^{n}$ that contains the true trajectory and the filter trajectory over the horizon of interest. On $\mathcal{X}$, the mappings $f_{\theta_{f}}$ and $h_{\theta_{h,i},i}$ are continuously differentiable and locally Lipschitz.
@@
    Bounded Jacobians on $\mathcal{X}$: There exist constants $M_{f},M_{h_{i}}>0$ such that
    $\sup_{\mathbf{x}\in\mathcal{X}}\|\nabla f_{\theta_{f}}(\mathbf{x})\|\leq M_{f},\quad\sup_{\mathbf{x}\in\mathcal{X}}\|\nabla h_{\theta_{h,i},i}(\mathbf{x})\|\leq M_{h_{i}}.$
    We do not require global bounds outside $\mathcal{X}$.
@@
    Noise and Initial Error Bounds: The process noise $\mathbf{w}_{k}$ and the measurement noise $\mathbf{v}_{k,i}$ are zero-mean and bounded in the mean-square sense. Moreover, there exists a constant $\delta>0$ such that the initial estimation error satisfies $\|\mathbf{e}_{0|0,i}\|<\delta$, for all nodes $i$.
  4. iv.
    Graph connectivity: The communication graph is (jointly) connected over the fusion horizon (e.g., $B$-connected).
  5. v.
    Innovation conditioning: There exists $s_{\min}>0$ such that the local innovation covariance satisfies $\mathbf{S}_{k,i}\succeq s_{\min}\mathbf{I}$ for all $k$ in the operating region (ensured in practice by $\mathbf{R}_{i}\succ 0$ and measurement excitation).
The Lipschitz/bounded-Jacobian conditions are enforceable during training using spectral-norm constraints, gradient penalties, or weight decay; empirical verification can be done by sampling $\|\nabla f_{\theta}\|$ and $\|\nabla h_{\theta,i}\|$ over a validation grid. The bound on $\mathbf{S}_{k,i}$ is a standard detectability/conditioning requirement and can be supported by measurement whitening or adaptive covariance inflation. The inequality $\|\mathbf{I}-\mathbf{K}_{k,i}\mathbf{H}_{\theta,i,k}\|\leq\beta_{i}$ is not designed directly (since $\mathbf{K}_{k,i}$ is computed from the Riccati recursion), but in practice can be promoted by: (a) maintaining reasonable $\mathbf{R}_{i}$ (well-conditioned innovations), (b) damping the update $\hat{\mathbf{x}}\leftarrow\hat{\mathbf{x}}+\alpha\,\mathbf{K}\,\mathrm{innov}$ with $\alpha\in(0,1]$, and (c) using CI when correlations are uncertain. Finally, the mean-value linearization remainder can be lumped into an effective process noise, i.e., $\mathbf{Q}\leftarrow\mathbf{Q}+\tilde{\mathbf{Q}}$, when state jumps are large.
@@ III-B Local Error Dynamics
Under local linearization, the prediction step yields an a priori error dynamics approximated by
$\mathbf{e}_{k+1|k,i}\approx\mathbf{F}_{\theta}(\xi_{k})\,\mathbf{e}_{k|k,i}+\mathbf{w}_{k},$
where $\xi_{k}$ lies on the line segment between $\mathbf{x}_{k}$ and $\hat{\mathbf{x}}_{k|k,i}$. Similarly, the measurement update gives
$\mathbf{e}_{k+1|k+1,i}\approx\Bigl(\mathbf{I}-\mathbf{K}_{k+1,i}\,\mathbf{H}_{\theta,i,k+1}\Bigr)\,\mathbf{e}_{k+1|k,i}-\mathbf{K}_{k+1,i}\,\mathbf{v}_{k+1,i}.$
The distributed fusion step is assumed to combine local estimates (e.g., via an information form aggregation) without degrading these contraction properties.
@@ III-C Sufficient Conditions for Stability
Theorem 1 (Local mean-square boundedness of NDKF)
Under Assumptions (i)–(v), suppose the true and estimated trajectories remain in $\mathcal{X}$ and that for each node $i$ there exists a constant $\beta_{i}\geq 0$ satisfying
$\|\mathbf{I}-\mathbf{K}_{k,i}\mathbf{H}_{\theta_{h,i},i,k}\|\leq\beta_{i},\;\forall\,k,\;\text{with}\;\gamma_{i}:=M_{f}\,\beta_{i}<1.$ (17)
Then the local estimation error satisfies
$\mathbb{E}\!\left[\|\mathbf{e}_{k|k,i}\|^{2}\right]\leq C_{i}\,\gamma_{i}^{2k}\,\|\mathbf{e}_{0|0,i}\|^{2}+\frac{D_{i}}{1-\gamma_{i}^{2}},$ (18)
for some constants $C_{i},D_{i}>0$ determined by the noise levels and the linearization remainder on $\mathcal{X}$. Hence NDKF is locally exponentially stable in mean square and converges to a bounded noise-dependent neighborhood.
Remark 2 (Interpretation of the assumptions)
The bounded-Jacobian condition is a local regularity assumption on the operating region, not a consequence of nor a substitute for global stability. It is practically checkable by spectral-normalization bounds and by sampling Jacobian norms on a validation grid. When the learned model error increases, its effect can be absorbed conservatively through covariance inflation, which preserves consistency at the cost of larger posterior uncertainty. Assumptions (iv)–(v) ensure, respectively, network connectivity for distributed fusion and well-conditioned innovations; they are not required for the local contraction bound but are used to preserve performance under distributed operation.
@@
We analyze the error propagation for a single node $i$ and then consider the fusion step. Let $\mathbf{e}_{k|k,i}=\mathbf{x}_{k}-\hat{\mathbf{x}}_{k|k,i}$ denote the local estimation error at time $k$, and assume that the local linearization is valid for sufficiently small errors.
@@
and the predicted state is given by
@@
By the mean value theorem, there exists a point $\xi_{k}$ between $\mathbf{x}_{k}$ and $\hat{\mathbf{x}}_{k|k,i}$ such that
@@
where $\mathbf{F}_{\theta}(\xi_{k})$ is the Jacobian of $f_{\theta}$ at $\xi_{k}$. By assumption,
$\|\mathbf{F}_{\theta}(\xi_{k})\|\leq M_{f}.$ (22)
Thus, the a priori error satisfies
$\|\mathbf{e}_{k+1|k,i}\|\leq M_{f}\,\|\mathbf{e}_{k|k,i}\|+\|\mathbf{w}_{k}\|.$ (23)
Remark 3 (Remainder term)
The mean-value form hides a higher-order remainder; under the Lipschitz-Jacobian assumption this remainder is bounded by $O(\|\mathbf{e}_{k|k,i}\|^{2})$ and can be conservatively absorbed into an effective process noise, which justifies the inequalities (23)–(28).
ii. Measurement Update. At the measurement update, the filter receives
@@
Subtracting the above from the true state (19), linearizing $h_{\theta,i}(\cdot)$ (with Jacobian $\mathbf{H}_{\theta,i,k+1}$ satisfying $\|\mathbf{H}_{\theta,i,k+1}\|\leq M_{h_{i}}$), and applying assumptions, yields
@@
By assumption, we have $\|\mathbf{I}-\mathbf{K}_{k+1,i}\mathbf{H}_{\theta,i,k+1}\|\leq\beta_{i}$, resulting in
$\|\mathbf{e}_{k+1|k+1,i}\|\leq\beta_{i}\left(M_{f}\|\mathbf{e}_{k|k,i}\|+\|\mathbf{w}_{k}\|\right)+\|\mathbf{K}_{k+1,i}\|\|\mathbf{v}_{k+1,i}\|.$ (28)
Defining $\gamma_{i}=M_{f}\,\beta_{i}<1$, we have the recursion:
$\|\mathbf{e}_{k+1|k+1,i}\|\leq\gamma_{i}\,\|\mathbf{e}_{k|k,i}\|+\beta_{i}\,\|\mathbf{w}_{k}\|+\|\mathbf{K}_{k+1,i}\|\,\|\mathbf{v}_{k+1,i}\|.$ (29)
iii. Induction. Iterating (29) from $k=0$ to $N$, we have
$\|\mathbf{e}_{N|N,i}\|\leq\gamma_{i}^{N}\,\|\mathbf{e}_{0|0,i}\|\\ +\sum_{j=0}^{N-1}\gamma_{i}^{N-1-j}\left(\beta_{i}\,\|\mathbf{w}_{j}\|+\|\mathbf{K}_{j+1,i}\|\|\mathbf{v}_{j+1,i}\|\right).$ (30)
Since $\gamma_{i}<1$, the term $\gamma_{i}^{N}\,\|\mathbf{e}_{0|0,i}\|$ decays exponentially, and under the bounded noise assumptions the summation converges. Hence, $\|\mathbf{e}_{N|N,i}\|$ converges exponentially fast to a bounded region determined by the noise covariances.
iv. Distributed Fusion Step. After local updates, each node $i$ fuses its estimate with neighbors:
@@
with weights $w_{ij}\geq 0$ and $\sum_{j\in\mathcal{N}_{i}}w_{ij}=1$. By convexity,
$\|\mathbf{e}^{(\mathrm{fusion})}_{k|k,i}\|\leq\max_{j\in\mathcal{N}_{i}}\|\mathbf{e}_{k|k,j}\|.$ (32)
Thus, the fusion step preserves the contraction properties established above.
Collectively, there exist constants $C,C^{\prime}>0$ such that
$\mathbb{E}\Bigl[\|\mathbf{e}_{k|k,i}\|^{2}\Bigr]\leq C\,\gamma_{i}^{2k}\,\|\mathbf{e}_{0|0,i}\|^{2}+\frac{C^{\prime}}{1-\gamma_{i}^{2}}.$ (33)
This establishes the exponential mean-square stability of the NDKF under the stated conditions. ∎
III-D Practical Proxies and Optimality
Instead of verifying $\|\mathbf{I}-\mathbf{K}\mathbf{H}\|\leq\beta$ directly, practitioners can monitor: (i) innovation conditioning ($\lambda_{\min}(\mathbf{S}_{k,i})\geq s_{\min}$), (ii) bounded covariance ($\operatorname{tr}(\mathbf{P}_{k|k})\leq p_{\max}$), and (iii) Jacobian norms ($\|\mathbf{F}_{\theta}(\cdot)\|\leq M_{f}$, $\|\mathbf{H}_{\theta,i}(\cdot)\|\leq M_{h_{i}}$) induced by training constraints. These proxies imply a contraction margin in the linearization region and are straightforward to track online.
With exact models, Gaussian noise, and valid local linearization, NDKF reduces locally to a distributed information filter/EKF and is locally MMSE on the linearization region. With learned and therefore imperfect models, NDKF is generally suboptimal; in that regime, the more relevant goal is consistency, which can be promoted through covariance inflation or covariance-intersection fusion.
@@ IV Results and Discussion
In this section, we present empirical evaluations of the proposed NDKF on a 2D system monitored by four distributed sensor nodes, and compare its performance to an Extended Kalman Filter (EKF) baseline.
@@ IV-A Experimental Setup
System Simulation: We consider a 2D state vector $\mathbf{x}_{k}=[p_{x},\,p_{y}]^{T},$ which evolves according to
$\mathbf{x}_{k+1}=\mathbf{x}_{k}+\begin{bmatrix}0.05\cos\bigl(k/10\bigr),0.05\sin\bigl(k/10\bigr)\end{bmatrix}^{\top}+\boldsymbol{\omega}_{k},$ (34)
@@
Neural Network Configuration and Training Process: To train the neural networks used in the NDKF, we adopt an offline learning strategy. First, we collect datasets consisting of state transition pairs $\{(\mathbf{x}_{k}^{(j)},\mathbf{x}_{k+1}^{(j)})\}_{j=1}^{M}$ for the dynamics network and measurement pairs $\{(\mathbf{x}_{k}^{(j)},\mathbf{y}_{k,i}^{(j)})\}_{j=1}^{M_{i}}$ for each sensor node’s measurement network. These datasets are split into training and validation sets. The dynamics network $f_{\theta}(\cdot)$ is implemented as a deep fully connected network with three hidden layers (128 neurons each), with batch normalization and dropout (0.2) to prevent overfitting. It is trained for 3000 epochs using the Adam optimizer with an initial learning rate of 0.001 and a learning rate scheduler. Similarly, each measurement network $h_{\theta,i}(\cdot)$ is implemented with two hidden layers (32 neurons each) and trained for 1000 epochs using analogous settings. The training objective minimizes the mean squared error loss functions in (3) and (4). Once convergence is achieved, the learned parameters are fixed and subsequently used in the online filtering process.
@@ IV-B Experimental Evaluation
The improvement in Table I is primarily attributable to reduced model mismatch, not to a different Kalman algebra. In the benchmark, the EKF baseline uses mis-specified observation functions at Nodes 1 and 2, whereas NDKF learns these mappings directly from data. As a result, the NDKF innovations are better centered and the local posteriors fused across nodes are less biased. We therefore interpret the improvement as evidence that learned models can be advantageous when analytical process or measurement equations are incomplete or inaccurate.
@@
Fig. 2 shows a representative 2D trajectory comparing the true state, the local node estimates, and the fused state estimate obtained using the NDKF. The NDKF’s fused estimate closely tracks the true state, while the EKF baseline deviates considerably, as indicated by the numerical RMSE in Table I, due to its mis‑specified measurement models. Fig. 3 illustrates the measurement innovation residuals for Node 1. The low and stable residuals obtained with the NDKF, along with the numerical results reported in Table I, confirm its superior measurement update performance compared to the EKF baseline.
@@
An implementation of NDKF, including the example provided in this section, is available in the GitHub repository accompanying this paper at: github.com/sfarzan/NDKF